OpenStudy (anonymous):

Help with math please! 4x^2+12x+5 how would i do this equation?

terenzreignz (terenzreignz):

Not an 'equation' unless there's an equals-sign. Is it equal to zero?

terenzreignz (terenzreignz):

Regardless, you might have to try and factor it anyway...

OpenStudy (anonymous):

Sorry i need to factor it

OpenStudy (anonymous):

Here, we can use factoring decomposition; determine two values in which the product is 20 and the sum is 12. They are 2 and 10, so split 12x into 2x and 12x! 4x² + 2x + 10x + 5 = 0 Factor like terms.. 2x(2x + 1) + 5(2x + 1) = 0 Group the common factors of 2x + 1 and collect the terms.. (2x + 1)(2x + 5) = 0 Finally, by zero-product property: 2x + 1 = 0 and 2x + 5 = 0 2x = -1 and 2x = -5 x = {-½, -5/2}

terenzreignz (terenzreignz):

Must defer. I'm admittedly not good at teaching how to factor, it just happens... T.T

OpenStudy (anonymous):

Also 16x^2-16x-12

OpenStudy (anonymous):

read about quadratic equations.. solve it by discriminant... u will find dat it has no real solution.. not every quadratic has real solution